OpenStudy (anonymous):

simplify: square root of 18 + square root of 50 a. 4 square root of 17 b. 2 square root of 17 c. 8 square root of 2 d. 16 square root of 2

OpenStudy (maheshmeghwal9):

just factor & try to take pair of 2 inside squareroot.

OpenStudy (maheshmeghwal9):

pair of 2 same terms it may be constant or variable

OpenStudy (maheshmeghwal9):

just as \[\sqrt{8}=\sqrt{2 \times 2 \times 2}=2\sqrt{2}\]

OpenStudy (maheshmeghwal9):